Let a_{k} = 3^{k} + k - 2 for all k \geq 0.


Write down the values of a_{1}, a_{2} and a_{3}.


Write down the values of A(1), A(2) and A(3) defined by the recurrence relation: A(0) = -1, A(k) = 3A(k - 1) - 2k + 7, k \geq 1


Show that A(k) = a_{k} is a solution of the recurrence relation for all values of k \geq 1.


1
Expert's answer
2022-02-28T17:25:39-0500

Given, "a_{k} = 3^{k} + k - 2 ~\forall k \geq 0."

Then,

"\begin{aligned}
a_{1} &= 3^1+1-2=2\\
a_2&=3^2+2-2=9\\
a_3&=3^3+3-2=28
\end{aligned}".


Given the recurrence relation, "A(k) = 3A(k - 1) - 2k + 7, k \geq 1" and "A(0) = -1".

"\begin{aligned}
A(1) &= 3A(0) - 2*1+7= -3-2+7=2\\
A(2) &= 3A(1) - 2*2+7=3*2-4+7=9\\
A(3) &= 3A(2)-2*3+7= 3*9-6+7=28
\end{aligned}"


The recurrence relation, "A(k) = 3A(k - 1) - 2k + 7, k \geq 1" can be written as

"A(k) - 3A(k - 1) = - 2k + 7, k \geq 1" which is a first order non-homogeneous relation.

The characteristic equation is "a - 3 = 0". The root is "a=3". Hence the homogeneous solution is "A_h(k) = c\cdot3^k".


Since the R.H.S of the given relation is "-2k+7", the particular solution is of the form "d_0 + d_1 k".

Using this in the given relation we get,

"\begin{aligned}
(d_0 + d_1 k) - 3(d_0 + d_1 (k-1)) &= 7 - 2k\\
d_0 + d_1 k - 3d_0 - 3d_1k +3d_1 &= 7 - 2k\\
(-2d_0+3d_1)-2d_1k &= 7-2k
\end{aligned}"


Equating the corresponding coefficients, we get

"\begin{aligned}
-2d_0+3d_1&=7\qquad \qquad \qquad \qquad (1)\\
-2d_1 &= -2 \implies d_1 = 1\\
\end{aligned}\\
\text{Using this in (1), we get~} d_0 = -2"


Hence the particular solution is "A_p(k) = k - 2."

Therefore, the solution of the recurrence relation is "A(k) = c\cdot 3^k + k - 2".


Given, A(0) = -1. Using this, we get

"-1=A(0) = c+0-2\implies c = 1"

Hence, "A(k) = 3^k + k - 2". Thus we proved "a_k" is the solution of "A(k)".